# Identifiable learning of dissipative dynamics

### v1

@@ Abstract
Complex dissipative systems appear across science and engineering, from polymers and active matter to learning algorithms. These systems operate far from equilibrium, where energy dissipation and time irreversibility are key to their behavior, but are difficult to quantify from data. Learning accurate and interpretable models of such dynamics remains a major challenge: the models must be expressive enough to describe diverse processes, yet constrained enough to remain physically meaningful and mathematically identifiable. Here, we introduce I-OnsagerNet, a neural framework that learns dissipative stochastic dynamics directly from trajectories while ensuring both interpretability and uniqueness. I-OnsagerNet extends the Onsager principle to guarantee that the learned potential is obtained from the stationary density and that the drift decomposes cleanly into time-reversible and time-irreversible components, as dictated by the Helmholtz decomposition. Our approach enables us to calculate the entropy production and to quantify irreversibility, offering a principled way to detect and quantify deviations from equilibrium. Applications to polymer stretching in elongational flow and to stochastic gradient Langevin dynamics reveal new insights, including super-linear scaling of barrier heights and sub-linear scaling of entropy production rates with the strain rate, and the suppression of irreversibility with increasing batch size. I-OnsagerNet thus establishes a general, data-driven framework for discovering and interpreting non-equilibrium dynamics.
@@ 1 Introduction
Complex dissipative dynamical systems are ubiquitous in both natural phenomena and engineering applications [18, 59, 55, 62, 10, 75], ranging from active matter and polymer physics to stochastic optimization algorithms in deep learning. While many foundational phenomena near or at equilibrium admit clear, mechanistically transparent descriptions, non-equilibrium systems lack a universally accepted canonical form and are instead described through empirical laws. Their defining features—energy dissipation, probability currents, and time irreversibility—are central to system function and performance [6, 38, 66, 2], yet they are difficult to quantify directly from observations and even harder to encode reliably in mathematical models. Building dynamical models that both predict accurately across conditions and reveal the organizing principles of non-equilibrium dynamics therefore remains a central challenge.
@@
Recent advances in artificial intelligence and machine learning have enabled powerful data-driven modeling of dynamical processes [7, 8, 12, 11, 41, 53, 29, 30, 28]. A particularly promising direction involves guiding neural networks with physical priors that constrain the hypothesis space, leading to predictive models that are both accurate and interpretable. Different from methods based on (sparse) identification from a dictionary [9, 8], this approach defines the hypothesis space using a continuous functional template that prescribes the intrinsic structure of admissible dynamics. Notable efforts in this direction include models based on Hamiltonian dynamics for conservative systems [3, 32, 81], and those built upon the Onsager principle or the GENERIC framework for dissipative, non-equilibrium dynamics [13, 36, 46, 77, 78].
@@
However, when such approaches are implemented, a fundamental dilemma arises. On the one hand, the hypothesis space must be expressive enough to capture a wide range of dissipative dynamics. On the other hand, it must be sufficiently constrained to guaranty model identifiability—ensuring, for example, that the learned energy function is unique up to an additive constant. Current methodologies for learning non-equilibrium dissipative dynamics do not satisfy these competing criteria simultaneously.
@@
In this paper, we propose I-OnsagerNet, an identifiable extension of OnsagerNet [77, 13] for learning dissipative, non-equilibrium stochastic dynamics from trajectory data. The hallmark of I-OnsagerNet is its physical interpretability: the potential function admits a direct interpretation as the negative logarithm of the stationary density, and the drift naturally decomposes into time-reversible and time-irreversible parts. These properties enable direct access to irreversibility and entropy production rate (EPR) [39, 19], allowing the model to diagnose departures from equilibrium and to quantify the underlying drivers of non-equilibrium behavior directly from data. Most importantly, our method achieves mathematical identifiability without sacrificing generality. We prove that I-OnsagerNet can represent any stochastic differential equation with a unique invariant distribution—covering the full class of stable and ergodic dissipative systems—and that its components are uniquely determined. Collectively, these features establish I-OnsagerNet as a general and identifiable framework for learning dissipative dynamics.
@@
We demonstrate the capabilities of I-OnsagerNet by learning the dynamics of a 300-bead polymer chain stretching in elongational flow, a system with up to 900 degrees of freedom. Using coarse-grained representations learned from microscopic data, we show that I-OnsagerNet reliably reconstructs consistent potential energy landscapes and reveals that the barrier height increases super-linearly with the strain rate—an observation not previously reported. In addition, I-OnsagerNet quantifies local and global entropy production rates under varying external conditions, exhibiting a sub-linear scaling of EPR with strain rate and confirming that the system operates far from equilibrium. We further validate these findings with single-molecule experimental data.
@@
To further demonstrate its broad applicability, we also employ I-OnsagerNet to investigate stochastic gradient Langevin dynamics, a widely used algorithm in Bayesian inference and non-convex optimization [75]. We quantify how mini-batching introduces state-dependent noise that breaks detailed balance and drives the system out of equilibrium. Our analysis reveals that the EPR decreases with the batch size before plateauing at larger batches, indicating a suppression of non-equilibrium behavior as stochasticity is reduced. In addition, we find that the sampling error follows a consistent trend, suggesting that the EPR can serve as a physically-grounded diagnostic for the sampling quality. Together, these applications illustrate the versatility of I-OnsagerNet in uncovering and quantifying non-equilibrium behavior in complex stochastic processes across physically and algorithmically generated non-equilibrium systems.
Figure 1: Overview of the I-OnsagerNet. (Blue) Discrete trajectory data serves as training input, and the model is built upon neural networks. (Green) Through a designed parameterization and likelihood-based training on this trajectory data, the framework constructs the I-OnsagerNet dynamical model. (Orange) Mathematical analysis establishes that I-OnsagerNet possesses both universal approximation capabilities and unique representation properties. (Grey) The components of the model can be interpreted as thermodynamically meaningful elements, including the stationary density, the time-reversible drift, and the time-irreversible drift. (Purple) Leveraging this physical interpretability and the identifiability, I-OnsagerNet enables the computation of the energy landscape and the EPR.
2 OnsagerNet for Modeling Dissipative Systems
OnsagerNet is a framework for learning the dynamics of dissipative systems from trajectory data [77]. It extends the Onsager principle to non-equilibrium regimes by incorporating both dissipative and conservative components. In its stochastic version [13], the system dynamics are expressed as
$\dot{\mathbf{Z}}_{t}=-[M(\mathbf{Z}_{t})+W(\mathbf{Z}_{t})]\nabla V(\mathbf{Z}_{t})+\sigma(\mathbf{Z}_{t})\dot{\mathbf{B}}_{t},$ (1)
where for each state $z$, $M(z)$ is a symmetric positive semi-definite matrix, $W(z)$ is an antisymmetric matrix capturing conservative motion, $V(z)$ is a potential function, and $\sigma(z)$ is the diffusion matrix. All four components are parameterized by neural networks and learned from data. This structure-preserving design enforces thermodynamic consistency, accommodates non-equilibrium behavior, and ensures long-term stability, distinguishing OnsagerNet from unconstrained neural ODE approaches [12]. However, the framework suffers from a critical non-identifiability issue. The decomposition into $V$, $M$, and $W$ is not unique, as the dynamics in (1) are invariant under the transformation $(V,M,W)\to(\varphi(V),M/\varphi^{\prime},W/\varphi^{\prime})$ for any strictly increasing differentiable function $\varphi$. Consequently, applying the method to the same dataset can yield a multitude of valid but distinct potential functions. This is a general issue for existing neural models designed to learn dissipative dynamics, such as GFINNs [78] and PHNNs [21]. We also mention the recent work [35] in which statistical-Physics-Informed Neural Networks were introduced to discover thermodynamic structures of equilibrium dissipative systems from short-time particle simulations. This work also addresses the uniqueness issue, but in a different setting from the present work which focuses on non-equilibrium systems.
3 I-OnsagerNet Architecture
To address the challenge of non-identifiability, we propose an identifiable extension of OnsagerNet framework, which we call I-OnsagerNet in this paper. Its dynamics are governed by the following Itô stochastic differential equation (SDE):
@@ 3 I-OnsagerNet Architecture
We employ a data-driven, machine-learning approach to reconstruct the stochastic dynamics in (2). The unknown functions $W$, $V$, and $\sigma$ are each parameterized as trainable deep neural networks. Essential physical priors are encoded into the neural network, including the positive definiteness of $M=\sigma\sigma^{\top}/2$ and the integrability of $\rho=e^{-V}$, which ensures the existence of the invariant distribution. The model is trained on a dataset of multiple independent trajectories by maximizing the likelihood of the observed data. Once trained, identifiability allows for subsequent analysis of the dynamical properties of the system. A schematic of this computational framework is provided in Fig. 1, with details of the network architecture and learning algorithm provided in Methods.
I-OnsagerNet naturally incorporates finite-temperature conditions, where temperature governs the noise intensity and influences thermodynamic quantities. Herein, we introduce a parameter $T$ representing temperature and rescale $\sigma=\sqrt{T}\tilde{\sigma}$ and $M=T\tilde{M}$. We further re-normalize the potential $V$ and the matrix $W$ as $V=\tilde{V}/T$ and $W=T\tilde{W}$. Substituting these rescaled quantities into the original formulation yields the $T$-dependent form:
It is noted that $\tilde{W}$ and $\tilde{V}$ remain dependent on $T$. Formally taking the limit $T\rightarrow 0$ and assuming that $\tilde{W}$ and $\tilde{V}$ have well-defined limits, we recover the deterministic OnsagerNet dynamics [77] as the leading-order approximation of I-OnsagerNet, differing only in the specific choice of the antisymmetric matrix $W$.
In comparison to S-OnsagerNet [13] which also accounts for stochastic effects, I-OnsagerNet employs specific matrix functions $M$ and $W$, and includes their divergences as distinct additive terms. These choices are not ad-hoc modifications but are principled constructions, emerging from the combination of the generalized Onsager principle [77], the generalized Helmholtz decomposition [19], and the essentially Hamiltonian decomposition [26]. The resulting dynamical model provides an identifiable extension of S-OnsagerNet and enables the computation of the invariant distribution and the entropy production rate (EPR), as demonstrated theoretically and numerically below.
Next, we show that the potential $V$ corresponds to the stationary density of the system, and the I-OnsagerNet drift, originally interpreted as dissipative and conservative forces, admits an equivalent decomposition into time-reversible and time-irreversible components. This property directly enables computation of the EPR for the learned system, a crucial physical metric quantifying its departure from equilibrium.
3.1 Stationary Density
@@
The probability density $\rho_{t}$ of the solution $\mathbf{Z}_{t}$ satisfies the Fokker-Planck equation (FPE) [61]:
@@
When the system has reached a statistical steady state, so that $\rho_{t}=\rho$, then the above FPE reduces to its stationary form [61]:
@@
Although stationary densities for general SDEs are typically intractable, it is straightforward to verify that for the I-OnsagerNet-defined SDE the stationary density takes the explicit form (the Appendix):
$\rho(\mathbf{Z})=\frac{1}{\mathcal{Z}}{e^{-V(\mathbf{Z})}},\quad\mathcal{Z}=\int e^{-V(\mathbf{Z})}\,d\mathbf{Z}.$
3.2 Time-reversal and Entropy Production
For a general time-homogeneous Markov process $\mathbf{Z}_{t}$ on $t\in[0,T]$, its time-reversed process $\bar{\mathbf{Z}}_{t}$ is defined as $\bar{\mathbf{Z}}_{t}=\mathbf{Z}_{T-t}$. For I-OnsagerNet, the precise correspondence between its potential function and the invariant distribution enables the explicit identification of the time-reversed dynamics. Specifically, when the forward process $\mathbf{Z}_{t}$ is stationary with respect to $\rho$, its time-reversed process $\bar{\mathbf{Z}}_{t}$ is the solution to the SDE of the form (the Appendix[19, 39]:
@@ 3.2 Time-reversal and Entropy Production
In other words, the governing function for the time-reversed process is obtained by leaving the $M$-related part unchanged while flipping the sign of the $W$-related part. It follows that the full dynamics of I-OnsagerNet can be split into time-reversible and time-irreversible components, which correspond exactly to the contributions from the matrices $M$ and $W$, respectively.
@@
Leveraging the physical interpretations, we are now able to calculate the entropy production using I-OnsagerNet.
The global EPR, $\dot{S}_{\text{tot}}$, quantifies the irreversibility of a stochastic process as the degree to which its trajectories are statistically distinguishable from their time-reversed counterparts. For a time-homogeneous Markov process, it is defined as the Kullback-Leibler (KL) divergence per unit time between the forward path measure $\mathbf{P}$ and the time-reversed path measure $\bar{\mathbf{P}}$ [17, 38, 39, 45]:
@@
where $\tau>0$ is arbitrary and $\{\mathbf{Z}_{t}\}_{t\in[0,\tau]}$ denotes a stationary path of the process. This formulation requires the simulation of the time-reversal dynamics and requires the calculation of the KL divergence between probability measures on path space, which presents computational challenges [14, 15]. For SDEs, it is well known that the global EPR can be represented in terms of a quadratic form for the time-irreversible drift [19, 39]. This provides a powerful analytical tool that is particularly accessible for I-OnsagerNet, as its reverse-time dynamics construction yields a closed-form expression. Specifically, $\dot{S}_{\text{tot}}$ for I-OnsagerNet can be expressed as (the Appendix):
@@
where $f_{\text{irr}}$ is the time-irreversible drift given by (4) and $\rho$ is the stationary density. Numerically, the I-OnsagerNet is first learned from input trajectories. The learned stochastic dynamics are then simulated over long time horizons to generate a large set of samples that approximate the invariant distribution. The global EPR is subsequently estimated from these samples using Monte Carlo integration of (6). By combining the above steps, I-OnsagerNet offers a fully data-driven pipeline to estimate irreversibility from raw trajectory data.
(6) indicates that $\dot{S}_{\text{tot}}=0$ is necessary and sufficient for equilibrium. To spatially quantify the breakdown of time-reversal symmetry and the emergence of non-equilibrium dynamics, we consider the (local) total EPR defined as the integrand of $\dot{S}_{\text{tot}}$ [63, 65, 66]:
@@
Beyond characterizing irreversibility, the time-irreversible drift $f_{\text{irr}}$ is identical to the current velocity and can be used to compute other measures such as the system EPR. These connections are discussed in detail in the Appendix.
3.3 Mathematical Characterization
We now present the main theorem concerning the expressive capacity and uniqueness of I-OnsagerNet, thereby establishing its identifiability from a theoretical standpoint. Detailed proofs can be found in the Appendix.
@@ 3.3 Mathematical Characterization
where $g$ denotes the drift vector. We will work under the standard assumptions on the drift and diffusion coefficients that ensure that the SDE admits a unique (strong) solution and a unique invariant distribution with density $\rho$ [61].
@@
For any SDE of the form (7), there exists a banded antisymmetric matrix $W$ with bandwidth 1 and a potential $V$ satisfying $e^{-V}\in L^{1}(\mathbb{R}^{D})$ such that the drift $g$ admits the following decomposition:
@@
on the support of the stationary density $\{\mathbf{Z}|\rho(\mathbf{Z})>0\}$. Furthermore, the potential $V$ is given by $V=-\log\rho$ and is unique up to an additive constant; the resulting potential force $\nabla V$, the time-reversible drift given by (3), and the time-irreversible drift given by (4) are uniquely determined.
Although existing modeling approaches offer considerable flexibility, the I-OnsagerNet is universal: it can represent any SDE admitting a unique invariant distribution and thus capture the dynamics of any stable and ergodic dissipative SDE [^1] An SDE satisfying a formal dissipativity condition with an appropriate Lyapunov function is guaranteed to possess a unique invariant distribution [37].. Here, the functions $\sigma$, $W$ and $V$ can be drawn from any hypothesis class that satisfies the classical universal-approximation theorem of functions, for example, feedforward, residual, or convolutional networks. Under this parameterization, the universal-approximation result for the proposed I-OnsagerNet follows directly from Theorem 1 and the approximation guaranties of the chosen class.
The uniqueness statement in the above theorem provides us with the key identifiable learning capacity of I-OnsagerNet. When applied to a family of SDEs indexed by a parameter, this means that the learned energy landscapes and time-irreversible drifts are guaranteed to be theoretically consistent across all parameter values. This consistency, even with minor numerical errors, provides a robust foundation for meaningfully comparing learned dynamical structures (e.g., energy landscapes, time-irreversible drift) across the parameter space within the SDE family. This will be demonstrated in our numerical examples. As an important byproduct of our proposed methodology, I-OnsagerNet can directly learn the invariant distribution from trajectory data alone. Although existing methods [50, 51] are also capable of learning energy-like functions to approximate the invariant distribution, either from trajectory data or from known governing equations, they typically rely on auxiliary loss terms derived from the stationary Fokker-Planck equation to promote consistency. In contrast, our method directly derives the potential function from the structure of the SDE, with a rigorous guarantee that the normalized density $\rho=\frac{1}{\mathcal{Z}}e^{-V}$ is exactly stationary for the learned dynamics.
4 Results

### v2

@@ Abstract
Complex dissipative systems appear across science and engineering, from polymers and active matter to learning algorithms. These systems operate far from equilibrium, where energy dissipation and time irreversibility govern their behavior but are difficult to quantify from data. Here, we introduce a universal and identifiable neural framework that learns dissipative stochastic dynamics directly from trajectories while ensuring interpretability, expressiveness, and uniqueness. Our method identifies a unique energy landscape, separates reversible from irreversible motion, and allows direct computation of the entropy production, providing a principled measure of irreversibility and deviations from equilibrium. Applications to polymer stretching in elongational flow and to stochastic gradient Langevin dynamics reveal new insights, including super-linear scaling of barrier heights and sub-linear scaling of entropy production rates with the strain rate, and the suppression of irreversibility with increasing batch size. Our methodology thus establishes a general, data-driven framework for discovering and interpreting non-equilibrium dynamics.
@@ 1 Introduction
Complex dissipative dynamical systems are ubiquitous in both natural phenomena and engineering applications [chaudhari2018stochastic, cross1993pattern, marchetti2013hydrodynamics, nartallo2026nonequilibrium, ottinger2005beyond, perkins1997single, welling2011bayesian], ranging from active matter and polymer physics to stochastic optimization algorithms in deep learning. While many foundational phenomena near or at equilibrium admit clear, mechanistically transparent descriptions, non-equilibrium systems lack a universally accepted canonical form and are instead described through empirical laws. Their defining features—energy dissipation, probability currents, and time irreversibility—are central to system function and performance [battle2016broken, boffi2024deep, jarzynski2012equalities, seifert2012stochastic], yet they are difficult to quantify directly from observations and even harder to encode reliably in mathematical models. Building dynamical models that both predict accurately across conditions and reveal the organizing principles of non-equilibrium dynamics therefore remains a central challenge.
@@
Recent advances in artificial intelligence and machine learning have enabled powerful data-driven modeling of dynamical processes [brunton2022data, brunton2016discovering, chen2025due, chen2018neural, gaskin2024neural, gaskin2023neural, gaskin2024inferring, karniadakis2021physics, liu2021machine]. A particularly promising direction involves guiding neural networks with physical priors that constrain the hypothesis space, leading to predictive models that are both accurate and interpretable. Different from methods based on (sparse) identification from a dictionary [brunton2016discovering, champion2019data], this approach defines the hypothesis space using a continuous functional template that prescribes the intrinsic structure of admissible dynamics. Notable efforts in this direction include models based on Hamiltonian dynamics for conservative systems [bertalan2019learning, greydanus2019hamiltonian, zhu2023machine], and those built upon the Onsager principle [chen2024constructing, yu2021onsagernet] or the GENERIC framework [lee2021machine, zhang2022gfinns] for dissipative, non-equilibrium dynamics.
@@
However, when such approaches are implemented, a fundamental dilemma arises. On the one hand, the hypothesis space must be expressive enough to capture a wide range of dissipative dynamics. On the other hand, it must be sufficiently constrained to guaranty model identifiability—ensuring, for example, that the learned energy function is unique up to an additive constant. Without identifiability, the learned representations cannot be used to reliably infer system behavior. This is particularly critical when learning from a family of datasets, where uniquely defined quantities are essential for systematic comparison across parameter space. Current methodologies for learning non-equilibrium dissipative dynamics do not satisfy these competing criteria simultaneously, as detailed in Table 1.
Methods Dynamics Universal Identifiable Non-trivial invariant transformation
Onsager principle [doi2011onsager, onsager1931reciprocal] $\dot{\mathbf{Z}}_{t}=-M(\mathbf{Z}_{t})\nabla V(\mathbf{Z}_{t})$ $(V,M)\to(\varphi(V),M/\varphi^{\prime})$
OnsagerNets [chen2024constructing, yu2021onsagernet] $\dot{\mathbf{Z}}_{t}=-[M(\mathbf{Z}_{t})+W(\mathbf{Z}_{t})]\nabla V(\mathbf{Z}_{t})$ $(V,M,W)\to(\varphi(V),M/\varphi^{\prime},W/\varphi^{\prime})$
GENERIC  [lee2021machine, zhang2022gfinns] $\dot{\mathbf{Z}}_{t}=L(\mathbf{Z}_{t})\nabla E(\mathbf{Z}_{t})+M(\mathbf{Z}_{t})\nabla S(\mathbf{Z}_{t})$ $(E,S,L,M)\to(\varphi(E),\varphi(S),L/\varphi^{\prime},M/\varphi^{\prime})$
Stat-PINNs [huang_al_2025] $\dot{\mathbf{Z}}_{t}=-M(\mathbf{Z}_{t})\nabla V(\mathbf{Z}_{t})+\sqrt{2\varepsilon M(\mathbf{Z}_{t})}\dot{\mathbf{B}}_{t}$ None
Ours Eq. (1) None
Table 1: Illustration of common frameworks for learning dissipative dynamics. The dynamics is invariant under the corresponding transformation, where $\varphi$ is any strictly increasing differentiable function. This invariance indicates that applying the method to the same dataset may yield multiple valid but distinct potential functions, thereby leading to non-identifiability. We also mention the recent work [huang_al_2025] in which Stat-PINNs were introduced to discover thermodynamic structures of equilibrium dissipative systems from short-time particle simulations. This work also addresses the uniqueness issue, but the model is formulated to represent only equilibrium systems.
@@
We introduce Universal and Identifiable Dissipative Dynamics (UIDD), a neural framework for learning dissipative, non-equilibrium stochastic dynamics directly from trajectory data. The hallmark of UIDD is the physical interpretability of its identified components: we construct a potential function which admits a direct interpretation as the negative logarithm of the stationary density, and a drift function that naturally decomposes into time-reversible and time-irreversible parts. These properties enable direct access to irreversibility and entropy production rate (EPR) [da2023entropy, jiang2004mathematical], allowing the model to diagnose departures from equilibrium and to quantify the underlying drivers of non-equilibrium behavior directly from data. Most importantly, our method achieves mathematical identifiability without sacrificing generality. We prove that UIDD can represent any stochastic differential equation with a unique invariant distribution—covering the full class of stable and ergodic dissipative systems—and that its components are uniquely determined.
@@
We demonstrate the capabilities of UIDD by learning the dynamics of a 300-bead polymer chain stretching in elongational flow, a system with up to 900 degrees of freedom. Using coarse-grained representations learned from microscopic data, we show that UIDD reliably reconstructs consistent potential energy landscapes and reveals that the barrier height increases super-linearly with the strain rate. In addition, UIDD further quantifies both local and global EPR, uncovering a sub-linear scaling of EPR with strain rate and confirming that the system operates far from equilibrium. These findings were previously inaccessible due to non-identifiability. We further validate the non-equilibrium nature with single-molecule experimental data.
@@
To further demonstrate its broad applicability, we also employ UIDD to investigate stochastic gradient Langevin dynamics, a widely used algorithm in Bayesian inference and non-convex optimization [welling2011bayesian]. We quantify how mini-batching introduces state-dependent noise that breaks detailed balance and drives the system out of equilibrium. Our analysis reveals that the EPR decreases with the batch size before plateauing at larger batches, indicating a suppression of non-equilibrium behavior as stochasticity is reduced. In addition, we find that the sampling error follows a consistent trend, suggesting that the EPR can serve as a physically-grounded diagnostic for the sampling quality. Together, these applications illustrate the versatility of UIDD in uncovering and quantifying non-equilibrium behavior in complex stochastic processes across physically and algorithmically generated non-equilibrium systems.
2 Results
2.1 UIDD Architecture
Universal and Identifiable Dissipative Dynamics (UIDD) is a neural framework for learning dissipative dynamics from discrete trajectory data. Its key innovation is a universal and identifiable dynamical formulation whose evolution is defined by the following Itô stochastic differential equation (SDE):
@@ 2.1 UIDD Architecture
We employ a data-driven, machine-learning approach to reconstruct the stochastic dynamics in Eq. (1). The unknown functions $W$, $V$, and $\sigma$ are each parameterized as trainable deep neural networks. Essential physical priors are encoded into the neural network, including the positive definiteness of $M=\sigma\sigma^{\top}/2$ and the integrability of $\rho=e^{-V}$, which ensures the existence of the invariant distribution. The model is trained on a dataset of multiple independent trajectories by maximizing the likelihood of the observed data. Once trained, identifiability allows for subsequent analysis of the dynamical properties of the system. A schematic of this computational framework is provided in Fig. 1, with details of the network architecture and learning algorithm provided in Methods C.
Figure 1: Overview of the UIDD. (Blue) Discrete trajectory data serves as training input, and the model is built upon neural networks. (Green) Through a designed parameterization and likelihood-based training on this trajectory data, the framework constructs the UIDD dynamical model. (Orange) Mathematical analysis establishes that UIDD possesses both universal approximation capabilities and unique representation properties. (Grey) The components of the model can be interpreted as thermodynamically meaningful elements, including the stationary density, the time-reversible drift, and the time-irreversible drift. (Purple) Leveraging this physical interpretability and the identifiability, UIDD enables the computation of the energy landscape and the EPR.
A comparison with existing dissipative dynamics models [chen2024constructing, huang_al_2025, lee2021machine, yu2021onsagernet, zhang2022gfinns] is provided in Methods A. Rather than ad-hoc modifications, UIDD emerges from the combination of the generalized Onsager principle [yu2021onsagernet], the generalized Helmholtz decomposition [da2023entropy], and the essentially Hamiltonian decomposition [feng1995volume], see Methods B.3. In UIDD, the potential $V$ specifies the stationary density; the drift admits an equivalent decomposition into time-reversible and time-irreversible components, enabling computation of the entropy production rate (EPR); and the resulting framework yields a universal and identifiable extension of existing dissipative models.
Stationary Density
@@
The probability density $\rho_{t}$ of the solution $\mathbf{Z}_{t}$ satisfies the Fokker-Planck equation (FPE) [pavliotis2014stochastic]:
@@
When the system has reached a statistical steady state, so that $\rho_{t}=\rho$, then the above FPE reduces to its stationary form [pavliotis2014stochastic]:
@@
Stationary densities of general SDEs are typically intractable. However, for the UIDD, the stationary density admits a closed-form expression (Methods B.1):
$\rho(\mathbf{Z})=\frac{1}{\mathcal{Z}}e^{-V(\mathbf{Z})},\qquad\mathcal{Z}=\int e^{-V(\mathbf{Z})}\mathop{}\!\mathrm{d}\mathbf{Z}.$
Time-reversal and Entropy Production
For a general time-homogeneous Markov process $\mathbf{Z}_{t}$ on $t\in[0,T]$, its time-reversed process $\bar{\mathbf{Z}}_{t}$ is defined as $\bar{\mathbf{Z}}_{t}=\mathbf{Z}_{T-t}$. For UIDD, the structured dynamics enables the explicit identification of the time-reversed dynamics. Specifically, when the forward process $\mathbf{Z}_{t}$ is stationary with respect to $\rho$, its time-reversed process $\bar{\mathbf{Z}}_{t}$ is the solution to the SDE of the form (Methods B.2[da2023entropy, jiang2004mathematical]:
@@
In other words, the governing function for the time-reversed process is obtained by leaving the $M$-related part unchanged while flipping the sign of the $W$-related part. It follows that the full dynamics of UIDD can be split into time-reversible and time-irreversible components, which correspond exactly to the contributions from the matrices $M$ and $W$, respectively.
@@
Leveraging the physical interpretations, we are now able to calculate the entropy production using UIDD.
The global EPR, $\dot{S}_{\text{tot}}$, quantifies the irreversibility of a stochastic process as the degree to which its trajectories are statistically distinguishable from their time-reversed counterparts. For a time-homogeneous Markov process, it is defined as the Kullback-Leibler (KL) divergence per unit time between the forward path measure $\mathbf{P}$ and the time-reversed path measure $\bar{\mathbf{P}}$ [crooks1999entropy, jarzynski2012equalities, jiang2004mathematical, lebowitz1999gallavotti]:
@@
where $\tau>0$ is arbitrary and $\{\mathbf{Z}_{t}\}_{t\in[0,\tau]}$ denotes a stationary path of the process. This formulation requires the simulation of the time-reversal dynamics and requires the calculation of the KL divergence between probability measures on path space, which presents computational challenges [chernyak2006path, chetrite2008fluctuation]. For SDEs, it is known that the global EPR can be represented in terms of a quadratic form for the time-irreversible drift [da2023entropy, jiang2004mathematical]. This provides a powerful analytical tool that is particularly accessible for UIDD, as its reverse-time dynamics construction yields a closed-form expression. Specifically, $\dot{S}_{\text{tot}}$ for UIDD can be expressed as (Methods B.2):
@@
where $f_{\text{irr}}$ is the time-irreversible drift given by Eq. (3) and $\rho$ is the stationary density. Numerically, the UIDD is first learned from input trajectories. The learned stochastic dynamics are then simulated over long time horizons to generate a large set of samples that approximate the invariant distribution. The global EPR is subsequently estimated from these samples using Monte Carlo integration of Eq. (5). By combining the above steps, UIDD offers a fully data-driven pipeline to estimate irreversibility from raw trajectory data.
$\dot{S}_{\text{tot}}=0$ is necessary and sufficient for equilibrium. To spatially quantify the breakdown of time-reversal symmetry and the emergence of non-equilibrium dynamics, we consider the (local) total EPR defined as the integrand of $\dot{S}_{\text{tot}}$ [pigolotti2017generic, seifert2005entropy, seifert2012stochastic]:
@@
Beyond characterizing irreversibility, the time-irreversible drift $f_{\text{irr}}$ is identical to the current velocity and can be used to compute other measures such as the system EPR. These connections are discussed in detail in Methods B.2.
Mathematical Characterization
We now present the main theorem concerning the expressive capacity and uniqueness of UIDD, thereby establishing its identifiability from a theoretical standpoint. Detailed proofs can be found in Methods B.3.
@@
where $g$ denotes the drift vector. We will work under the standard assumptions on the drift and diffusion coefficients that ensure that the SDE admits a unique (strong) solution and a unique invariant distribution with density $\rho$ [pavliotis2014stochastic].
@@
For any SDE of the form Eq. (6), there exists a banded antisymmetric matrix $W$ with bandwidth 1 and a potential $V$ satisfying $e^{-V}\in L^{1}(\mathbb{R}^{D})$ such that the drift $g$ admits the following decomposition:
@@
on the support of the stationary density $\{\mathbf{Z}|\rho(\mathbf{Z})>0\}$. Furthermore, the potential $V$ is given by $V=-\log\rho$ and is unique up to an additive constant; the resulting potential force $\nabla V$, the time-reversible drift given by Eq. (2), and the time-irreversible drift given by Eq. (3) are uniquely determined.
Although existing modeling approaches offer considerable flexibility, UIDD is theoretically universal: it can represent any SDE admitting a unique invariant distribution and thus capture the dynamics of any stable and ergodic dissipative SDE [^1] An SDE satisfying a formal dissipativity condition with an appropriate Lyapunov function is guaranteed to possess a unique invariant distribution [huang2015steady].. Here, the functions $\sigma$, $W$ and $V$ can be drawn from any hypothesis class that satisfies the classical universal-approximation theorem of functions, for example, feedforward, residual, or convolutional networks. Under this parameterization, the universal-approximation result for the proposed UIDD follows directly from Theorem 1 and the approximation guaranties of the chosen class.
The uniqueness statement in the above theorem provides us with the key identifiable learning capacity of UIDD. When applied to a family of SDEs indexed by a parameter, this means that the learned energy landscapes and time-irreversible drifts are guaranteed to be theoretically consistent across all parameter values. This consistency, even with minor numerical errors, provides a robust foundation for meaningfully comparing learned dynamical structures (e.g., energy landscapes, time-irreversible drift) across the parameter space within the SDE family. This will be demonstrated in our numerical examples. As an important byproduct of our proposed methodology, UIDD can directly learn the invariant distribution from trajectory data alone. Although existing methods [lin2022computing, lin2023computing] are also capable of learning energy-like functions to approximate the invariant distribution, either from trajectory data or from known governing equations, they typically rely on auxiliary loss terms derived from the stationary Fokker-Planck equation to promote consistency. In contrast, our method directly derives the potential function from the structure of the SDE, with a rigorous guarantee that the normalized density $\rho=\frac{1}{\mathcal{Z}}e^{-V}$ is exactly stationary for the learned dynamics.